OpenStudy (anonymous):

How to use the distributive property to eliminate the parentheses? -13(x squared-2x+4)=

OpenStudy (anonymous):

\[-13(x^2-2x+4)=-13\times x^2-13\times (-2x)+(-13)\times4\]is a start

OpenStudy (anonymous):

The distribute property, in one of the directions, says that a ( b + c ) = a * b + a * c. The distributive property holds on the field of real numbers, which is where your numbers are.

OpenStudy (anonymous):

a, b, c above are any real numbers.

OpenStudy (anonymous):

thank you very much!